-2/5(3x+1)=1/3(2-x)
Answer by user_dude2008(1862)

-2/5(3x+1)=1/3(2-x)


-6(3x+1)=5(2-x)


-18x-6=10-5x

-18x+5x=10+6

-13x=16

x=-16/13